[Brahms-l] Shift report 20040316 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 16 2004 - 00:01:05 EST
------------------------------------------------------------
Date: Tue Mar 16 00:01:04 2004, Shift: 16:00-24:00
Supervisor: CEJ, with: SN
Log book#15 pages 89 - 89

Shift Summary:
------------------------------------------------------------
1) Beam untill 18:00. MCR called at 16:45, the end of the store was used for a PHENIX scan - we couldn't take data there. Beam again from 18:35 to 23:10.

2) Pedestal runs 11052 (TPC) and 11053.
   FS@8degA3540, MRS@90degB2255
   11045; tr2: 12K, tr3: 15K 
   11058;      20K,      15K
   11059;      40K,      51K
   11061;      51K,      65K
   11062;      36K,      46K
   11063;      29K,      37K

3) Still problems with Run- and IO server.
   BBL tube 16 TDC is gone. ADC looks fine.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 16 00:01:14 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 16 2004 - 00:01:35 EST